# Examining Algorithmic Curation on Social Media: An Empirical Audit of Reddit’s r/popular Feed

### v1

@@ Abstract
Platforms are increasingly relying on algorithms to curate the content within users’ social media feeds. However, the growing prominence of proprietary, algorithmically curated feeds has concealed what factors influence the presentation of content on social media feeds and how that presentation affects user behavior. This lack of transparency can be detrimental to users, from reducing users’ agency over their content consumption to the propagation of misinformation and toxic content. To uncover details about how these feeds operate and influence user behavior, we conduct an empirical audit of Reddit’s algorithmically curated trending feed called r/popular. Using 10K r/popular posts collected by taking snapshots of the feed over 11 months, we find that the total number of comments and recent activity (commenting and voting) helped posts remain on r/popular longer and climb the feed. Using over 1.5M snapshots, we examine how differing ranks on r/popular correlated with engagement. More specifically, we find that posts below rank 80 showed a sharp decline in activity compared to posts above, and that posts at the top of r/popular had a higher proportion of undesired comments than those lower down. Our findings highlight that the order in which content is ranked can influence the levels and types of user engagement within algorithmically curated feeds. This relationship between algorithmic rank and engagement highlights the extent to which algorithms employed by social media platforms essentially determine which content is prioritized and which is not. We conclude by discussing how content creators, consumers, and moderators on social media platforms can benefit from empirical audits aimed at improving transparency in algorithmically curated feeds.
@@ 1 Introduction
Social media platforms are flooded with immense amounts of new content every day. This constant stream of content has led to the reliance on algorithms to curate the content in users’ social media feeds. Algorithmic curation is defined as the process of “organizing, selecting, and presenting subsets of a corpus of information for consumption” (Rader and Gray 2015). In recent times, the most prominent examples of algorithmic curation are on short-form video platforms like TikTok, Instagram Reels, and YouTube Shorts. However, less thought of are the “hot” and “popular” feeds that exist on multiple platforms (e.g., GitHub, Reddit, X/Twitter) that also use algorithms to curate what is “trending.”
@@
The r/popular feed111https://www.reddit.com/r/popular on Reddit is one of the default feeds that is available to all users, with and without an account. The posts on the r/popular feed come from nearly all consenting communities on the platform—however, there are some subreddits, particularly “not safe for work” (NSFW) communities, that are omitted. The r/popular feed is a ranked list (see Figure 1) where the rankings are based on a calculated “hot” score (likely based on recency, comment rate, and upvote score) as well as some measures for post and discussion quality.222www.reddit.com/r/changelog/comments/9n3ix9/rpopular˙is˙changing/ At a high level, the r/popular feed consolidates the most active posts on Reddit from numerous subreddits and serves as the front page to the platform.
@@
What factors affect how long a post stays on the r/popular feed (i.e., the post’s tenure on the feed)?
What factors affect the assigned rank/position of the post on the r/popular feed?
How does the post’s rank on the r/popular feed affect the engagement on the post?
@@
Summary of Findings
Through our analyses, we find that the total number of comments, along with recent commenting and voting activity, were the most predictive factors for a longer stay on the r/popular feed (RQ1) and for upward movement on the feed (RQ2). We also find that undesired comments, i.e., toxic and moderator-removed comments, were also predictive of a longer stay and upward movement, but to a lesser degree. Regarding how ranking on r/popular affects engagement (RQ3), we find that posts which were higher on the feed received comments at a higher rate, as well as a greater proportion of undesired comments.
@@ 2 Related Work
Popularity on social media has been studied before; however, these studies often focus on how users (DeVito 2022), communities (Maldeniya et al. 2020; Chan et al. 2024), or topics (Schlessinger et al. 2023) change after becoming viral. For example, Gurjar et al. (2022) found that after users became viral, they increased their posting frequency and changed their posts to be similar to the post that made them viral. In general, the focus of these studies is the impact of popularity on (groups of) users. However, missing from prior work are investigations into why content or users became viral in the first place. To address this gap, we examine the factors that influence the systems that determine what is viral on social media, specifically the r/popular feed on Reddit. It is important to understand how these algorithmic curation systems work because they are pervasive on social media and the popularity they induce can be a double-edged sword, particularly for those of marginalized identities (DeVito 2022). Additionally, these systems can also propagate inflammatory content as noted by mainstream media in recent times (O’Sullivan 2019; Katherine J. Wu 2019; Faddoul, Chaslot, and Farid 2020).
@@
According to Rader and Gray (2015), algorithmic curation is the process of “organizing, selecting, and presenting subsets of a corpus of information for consumptions.” Within this definition, we are particularly focused on the organizing and presenting aspect because Reddit algorithmically ranks posts on r/popular. However, there exist other definitions (Cotter, Cho, and Rader 2017; Eckles 2022; He et al. 2023). Algorithmic ranking is only one key mechanism in which algorithmic curation systems can exert power (Diakopoulos 2014). By ranking content, specifically on social media feeds, these algorithms essentially have the power to determine what is important by prioritizing content or users over one another. To show how impactful algorithmic ranking can be, Joachims et al. (2017) examined users’ clickthrough behavior on Google’s result page and found that participants’ trust in Google’s retrieval/ranking function led them to click on highly ranked links regardless of their quality or relevance to the query. Additionally, Salganik, Dodds, and Watts (2006) found that ranking songs in music markets by total downloads produced more unpredictability and inequality compared to groups who had songs ranked randomly. We extend this line of work by examining the influence of algorithmic ranking on social media trending feeds.
@@ 3 Data
For tractability, we randomly sampled 10,000 posts from the 134,661 posts observed. The resulting sample contains 694 subreddits and 1,548,266 snapshots. For the rest of the paper, our findings are based on this representative sample of 10,000 posts and their respective snapshots. From the sample, we found that posts stay on r/popular for about 164 snapshots on average (, ). Additionally, posts from the sample, on average, stay on r/popular’s top 100 for 6.1 hours (, ). Along with these snapshots, we used Pushshift (Baumgartner et al. 2020) to obtain the comments for each post.
@@
In recent years, news outlets have suggested that social media platforms are intentionally promoting antisocial content through algorithmic prioritization to drive greater user engagement (Katherine J. Wu 2019; O’Sullivan 2019). To investigate whether antisocial behavior has any interactions with algorithmic curation, specifically on r/popular, we employed Almerekhi, Kwak, and Jansen (2022)’s fine-tuned BERT model, which assigns three toxicity-related scores for each comment: non_toxic, slightly_toxic, and highly_toxic. Each score ranges from 0 to 1. If a comment exceeded a score of 0.5—a threshold used in prior work (Bao et al. 2021)—for either slightly_toxic or highly_toxic, then we labeled it as toxic. Almerekhi, Kwak, and Jansen (2022) fine-tuned the model using r/AskReddit comments and showed its generalizability to 99 other large subreddits. Since the subreddits that appear on r/popular are mostly large, we claim that using this model is well suited for identifying toxic comments made within our r/popular posts. Additionally, some comments were removed, presumably by a moderator or bot, before they could be archived by Pushshift. To better capture antisocial behavior, we combine comments that contained “[removed]” with the ones flagged by the BERT model under an umbrella term: “undesired comments.”
@@
In this section, we describe the features we used for our regression models and provide a brief justification for inclusion. This feature list applies to posts captured at a specific snapshot/time. Thus, the features are measured at a specific time, e.g., the number of comments at a particular snapshot.
@@
We included recent activity because there is a “hot” calculation2 that likely includes comment and vote velocity, i.e., how many comments are coming in currently. A 10-minute window was selected because posts stay at a rank for an average 7.5 of minutes before moving onto another (, ). Calculating the proportion of undesired comments helps us test whether Reddit’s ranking algorithm is in any way influenced by the presence of undesired activity.
@@
Lastly, communities are integral to Reddit, which is why we included their size as a control variable in the models.
Table 1 provides the geometric mean and standard deviations for each feature. We used the geometric versions of these measures because the variables are log-transformed in our regression analyses. Additionally, the geometric standard deviations are used to inform the units we used to scale the regression coefficients which are also shown in Table 1. The consistency to use 2x for all units except the proportion of upvotes is to assist with interpretability.
@@ 4 RQ1: Tenure on r/popular
To recall, RQ1 asks what factors are associated with how long a post stays on r/popular (i.e., tenure). To conduct this analysis, we built logistic regression models that estimate whether the post continues to exist on the feed in the next snapshot—approximately 2 minutes later. This section describes the model and its findings in further detail.
Model 1: Logistic Regression
@@ Results
Impact of Overall Engagement. We found the strongest factors that increased the odds of staying near the top of the feed were the number of total comments, recent comments, and recent votes. We also found age (i.e., time since post creation) to be a strong factor that decreased the odds of staying near the top. These are consistent across all rank intervals and are natural for an engagement-based popular feed.
We also observed that increasing the score on a post corresponded to a drastic decrease in the odds of staying on the feed. We further discuss how this counterintuitive result may indicate the correlation between a post’s score and some unobserved influence in Section 8.
Impact of Undesired Activity. We found that increasing the proportion of undesired comments and the number of subscribers had a comparatively small but significant effect on increasing the odds that a post stays within the top 50 and top 25, with a stronger effect in the top 10. While these effects do not necessarily indicate that the ranking algorithm is intentionally designed to promote antisocial content to drive more user engagement, these results fail to rule it out. These effects also hint at an interaction between undesired comments and the rank of a post, and that the effects of undesired comments are strongest for posts at the top of the feed.

### v3

@@ Abstract
Platforms are increasingly relying on algorithms to curate the content within users’ social media feeds. However, the growing prominence of proprietary, algorithmically curated feeds has concealed what factors influence the presentation of content on social media feeds and how that presentation affects user behavior. This lack of transparency can be detrimental to users, from reducing users’ agency over their content consumption to the propagation of misinformation and toxic content. To uncover details about how these feeds operate and influence user behavior, we conduct an empirical audit of Reddit’s algorithmically curated trending feed called r/popular. Using 10K r/popular posts collected by taking snapshots of the feed over 11 months, we find that recent comments help a post remain on r/popular longer and climb the feed. We also find that posts below rank 80 correspond to a sharp decline in activity compared to posts above. When examining the effects of having a higher proportion of undesired behavior—i.e., moderator-removed and toxic comments—we find no significant evidence that it helps posts stay on r/popular for longer. Although posts closer to the top receive more undesired comments, we find this increase to coincide with a broader increase in overall engagement—rather than indicating a disproportionate effect on undesired activity. The relationships between algorithmic rank and engagement highlight the extent to which algorithms employed by social media platforms essentially determine which content is prioritized and which is not. We conclude by discussing how content creators, consumers, and moderators on social media platforms can benefit from empirical audits aimed at improving transparency in algorithmically curated feeds.
@@ 1 Introduction
Social media platforms are flooded with immense amounts of new content every day. This constant stream of content has led to the reliance on algorithms to curate the content in users’ social media feeds. Algorithmic curation is defined as the process of “organizing, selecting, and presenting subsets of a corpus of information for consumption” (Rader and Gray 2015). In recent times, the most prominent examples of algorithmic curation are on short-form video platforms like TikTok, YouTube Shorts, and Instagram Reels. However, less thought of are the “hot” and “popular” feeds on multiple platforms (e.g., GitHub, Reddit, X/Twitter) that also use algorithms to curate what is “trending” (Gillespie 2016).
@@
Before describing r/popular in further detail, we emphasize the importance of transparency in algorithmic curation—particularly ranking. Ranking shapes online behavior (Salganik, Dodds, and Watts 2006; Piccardi et al. 2024) and is often interpreted as a signal of quality (Joachims et al. 2017). This gives platforms significant influence over what content receives attention and how users engage with it (Poudel and Weninger 2024), with downstream effects on political beliefs (Epstein and Robertson 2015; Kulshrestha et al. 2017) and the ability to assess truth online (Pogacar et al. 2017). We extend research on algorithmic transparency by examining what drives and is driven by content ranking in social media feeds.
@@
The r/popular feed111https://reddit.com/r/popular on Reddit is one of the default feeds that is available to all users, with and without an account. The posts on the r/popular feed come from nearly all consenting communities on the platform—however, there are some subreddits, particularly “not safe for work” (NSFW) communities, that are omitted. The r/popular feed is a ranked list (see Figure 1) where the rankings are based on a calculated “hot” score (likely based on recency, comment rate, and upvote score) as well as some measures for post and discussion quality.222https://reddit.com/9n3ix9 At a high level, the r/popular feed consolidates the most active posts on Reddit from numerous subreddits and serves as the front page to the platform.
@@
What factors are associated with how long a post stays on the r/popular feed (i.e., the post’s tenure on the feed)?
What factors are associated with the assigned rank/position of the post on the r/popular feed?
How is the post’s rank on the r/popular feed associated with the engagement on the post?
@@
Through our analyses, we find that the total number of comments, along with recent commenting and voting activity, were the most predictive factors for a longer stay on the r/popular feed (RQ1) and for upward movement on the feed (RQ2). We also find that undesired comments (i.e., toxic and moderator-removed comments) were also predictive of a longer stay and upward movement, but to a lesser degree. Regarding how ranking on r/popular affects engagement (RQ3), we find that posts which were higher on the feed received comments at a higher rate, as well as a greater proportion of undesired comments.
We employ multiple regression models to obtain relationships between post engagement and rank/position or movement on the r/popular feed. To complement these correlational insights, we employ a causal inference approach drawing on the potential outcomes framework (Rubin 2005) to assess causal relationships (detailed in Section 4). This analysis provides evidence that increased commenting activity helps posts remain on r/popular longer (RQ1) and move up the feed (RQ2). We also find that upward movement on the feed leads to more commenting engagement (RQ3), including both overall and undesired engagement. However, we observe no disproportionate effect of rank on undesired engagement specifically.
@@ 2 Related Work
Popularity on social media has been studied before; however, these studies often focus on how users (DeVito 2022), communities (Maldeniya et al. 2020; Chan et al. 2024), or topics (Schlessinger et al. 2023) change after becoming viral. For example, Gurjar et al. (2022) found that after users became viral, they increased their posting frequency and changed their posts to be similar to the post that made them viral. In general, the focus of these studies is the impact of popularity on (groups of) users. However, missing from prior work are investigations into why content or users became viral in the first place. To address this gap, we examine the factors that influence the systems that determine what is viral on social media, specifically the r/popular feed on Reddit. It is important to understand how these algorithmic curation systems work because they are pervasive on social media and the popularity they induce can be a double-edged sword, particularly for those of marginalized identities (DeVito 2022). Additionally, these systems can also propagate inflammatory content, as noted by mainstream media in recent times (O’Sullivan 2019; Wu 2019).
@@
According to Rader and Gray (2015), algorithmic curation is the process of “organizing, selecting, and presenting subsets of a corpus of information for consumptions.” Within this definition, we are particularly focused on the organizing and presenting aspect because Reddit algorithmically ranks posts on r/popular. However, there exist other definitions (Cotter, Cho, and Rader 2017; Eckles 2022; He et al. 2023). Algorithmic ranking is only one key mechanism in which algorithmic curation systems can exert power (Diakopoulos 2013). By ranking content, specifically on social media feeds, these algorithms essentially have the power to determine what is important by prioritizing content or users over one another. To show how impactful algorithmic ranking can be, Joachims et al. (2017) examined users’ clickthrough behavior on Google’s result page and found that participants’ trust in Google’s retrieval/ranking function led them to click on highly ranked links regardless of their quality or relevance to the query. Additionally, Salganik, Dodds, and Watts (2006) found that ranking songs in music markets by total downloads produced more unpredictability and inequality compared to groups who had songs ranked randomly. We extend this line of work by examining the influence of algorithmic ranking on social media trending feeds.
@@ 3 Data
For tractability, we randomly sampled 10,000 posts from the 134,661 posts observed. The resulting sample contains 694 subreddits and 1,548,266 snapshots. For the rest of the paper, our findings are based on this representative sample of 10,000 posts and their respective snapshots. From the sample, we found that posts stay on r/popular for about 164 snapshots on average (, ). Additionally, posts from the sample, on average, stay on r/popular’s top 100 for 6.1 hours (, ). Along with these snapshots, we use Pushshift (Baumgartner et al. 2020) to obtain the comments for each post.
@@
In recent years, news outlets have suggested that social media platforms are intentionally promoting antisocial content through algorithmic prioritization to drive greater user engagement (Wu 2019; O’Sullivan 2019). To investigate whether antisocial behavior has any interactions with algorithmic curation, specifically on r/popular, we employ Almerekhi, Kwak, and Jansen (2022)’s fine-tuned BERT model, which assigns three toxicity-related scores for each comment: non_toxic, slightly_toxic, and highly_toxic. Each score ranges from 0 to 1. If a comment exceeds a score of 0.5—a threshold used in prior work (Bao et al. 2021)—for either slightly_toxic or highly_toxic, then we label it as toxic. Almerekhi, Kwak, and Jansen (2022) fine-tuned the model using r/AskReddit comments and showed its generalizability to 99 other large subreddits. Since the subreddits that appear on r/popular are mostly large, we claim that using this model is well suited for identifying toxic comments made within our r/popular posts. Additionally, some comments were removed, presumably by a moderator or bot, before they could be archived by Pushshift. To better capture antisocial behavior, we combine comments that contained “[removed]” with the ones flagged by the BERT model under an umbrella term: “undesired comments.”
@@
In this section, we describe the features we use for our regression models and provide a brief justification for inclusion. This feature list applies to posts captured at a specific snapshot/time. Thus, the features are measured at a specific time, e.g., the number of comments at a particular snapshot.
@@
We include recent activity because there is a “hot” calculation2 that likely includes comment and vote velocity, i.e., how many comments are coming in currently. We define recent activity as the activity in the past 10 minutes because posts stay at a rank for an average 7.5 of minutes before moving onto another (, ). Calculating the proportion of undesired comments helps us test whether Reddit’s ranking algorithm is in any way influenced by the presence of undesired activity.
@@
Lastly, communities are integral to Reddit, which is why we include their size as a control variable in the models.
Table 1 provides the geometric mean and standard deviations for each feature. We use the geometric versions of these measures because the variables are log-transformed in our regression analyses. Additionally, the geometric standard deviations are used to inform the units we use to scale the regression coefficients which are also shown in Table 1. The consistency to use 2x for all units except the proportion of upvotes is to assist with interpretability.
@@
4 Overarching Causal Inference Design
To complement our regression analysis, we adopt a causal inference approach drawing on the potential outcomes framework (Rubin 2005). Our goal is to examine the causal effects of a post receiving various “treatments” (e.g., moving up the feed) compared to what would have happened if the treatments were not administered (i.e., the “counterfactuals”). Although our work relies on observational data, the potential outcomes framework allows us to estimate the missing “counterfactual” for each treated post based on the outcomes of similar (i.e., matched), observed posts that were not treated. In other words, matching enables a quasi-experimental setup to estimate counterfactuals while accounting for numerous covariates.
We use this causal framework to conduct a time series analysis, matching on covariates measured in the snapshots leading up to the treatment. We employ stratified propensity score matching (Austin 2011) to identify comparable control snapshots for each treatment. Following matching, we apply generalized linear models (GLMs) to estimate the effects of the treatments on outcomes measured in the snapshots immediately following the treatment.
Treatment Definitions
The treatments in our causal analysis include changes in the number of recent comments (TC), the proportion of undesired comments during that time (TU), and rank movements on the r/popular feed (TR). We define treatment and control groups using consistent percentile-based thresholds, and we define weak and strong versions of each treatment.
For recent comments, the weak treatment (TC-W) is defined as receiving more comments in the past 10 minutes than the median ( comments), and the strong treatment (TC-S) is defined as receiving at least the 75th percentile ( comments). The control is defined as receiving at most the median ( comments).
For undesired comments, we define treatments based on the proportion of undesired comments received in the past 10 minutes. We use the proportion of undesired comments to define our treatment instead of the number of undesired comments, as the latter would be confounded by the overall number of recent comments. The weak treatment (TU-W) is defined as receiving a proportion of undesired comments greater than the median (), the strong treatment (TU-S) is greater than or equal to the 75th percentile (), and the control is less than or equal to the median ().
Finally, for rank movement, the weak treatment (TR-W) is defined as moving at least one rank up the feed, and the strong treatment (TR-S) is defined as moving up by at least five ranks (75th percentile for rank movement). If the post does not climb the feed—i.e., either stays in the same position or falls—then it is considered a control sample.
These cutoffs are summarized in Appendix B, which also includes information regarding filtering. For each treatment, we estimate its effect on outcomes of tenure (RQ1), rank movement (RQ2), and engagement (RQ3).
Stratified Propensity Score Matching
For a quasi-experimental design, we conduct matching to find comparable treated and control groups. A propensity score model matches snapshots based on their likelihood of receiving the treatment, quantified as propensity scores. In particular, we conduct stratified propensity score matching—which balances the bias-variance tradeoff of comparing too biased (one-to-one matched) or too variant (unmatched) samples (Kiciman, Counts, and Gasser 2018). This approach distributes groups with similar propensity scores into strata, and is a well-established method adopted in prior work (Saha et al. 2021; Verma et al. 2022; Yuan et al. 2023; Lambert, Saha, and Chandrasekharan 2025).
For each treatment, we train an AdaBoost classifier to estimate propensity scores using the following covariates: (1) post age, (2) title text (top 200 bi-grams), (3) rank, (4) time at rank, (5) number of comments, (6) number of undesired comments, (7) overall comment rate, and (8) undesired comment rate. Snapshots are then sorted by their propensity scores and partitioned into equally sized strata, such that snapshots within a stratum are similarly likely to receive the treatment based on the covariates.
To ensure high-quality strata, we exclude any stratum with fewer than 200 treatment or control snapshots, as well as strata where the standardized mean difference (SMD) between treatment and control groups exceeds 0.3—a common threshold in prior work (Kampenes et al. 2007; Kiciman, Counts, and Gasser 2018). This matching approach approximates the presence of a counterfactual snapshot that was not treated, enabling us to assess causal relationships between treatments and outcomes. Further details on the number of strata pre- and post-filtering are provided in Appendix B.
Generalized Linear Models (GLMs)
After filtering for high-quality strata, we employ GLMs and use Bayesian inference to estimate whether the treatment had any effect on the outcome of interest. We use a hierarchical Bayesian model to pool effects across strata, where within-stratum effects are sampled from a global distribution whose mean represents the average treatment effect.
We present our findings with probability of direction () and region of practical equivalence () (full) as these are commonly used when communicating the strength and significance of findings in a Bayesian context (Makowski et al. 2019). The probability in can be interpreted as the probability that a significant effect does not exist, and can be interpreted as the probability that an effect is in a particular direction. To facilitate discussion of whether the effect size is large enough to be “significant,” we define our to be between -4.8% and 5%, and we consider effect sizes within the ROPE to be “not significant.” We also report the mean of the estimated effect size along with 95% credible interval.
@@ 5 RQ1: Tenure on r/popular
RQ1 asks what factors are associated with how long a post stays on r/popular (i.e., tenure). To conduct this analysis, we built logistic regression models to estimate whether the post continues to exist on the feed in the next snapshot—approximately 2 minutes later. Additionally, we fit a set of GLMs to assess causal links between commenting activity and tenure. This section describes the models and their findings in further detail.
Logistic Regression
@@
Causal Inference Analysis
We examined the causal relationship between commenting activity and post tenure by using a set of logistic regression models. We apply these models on our stratified data to estimate the effects of recent commenting activity (TC) and undesired activity (TU) on the odds of appearing within the top 50, 25, 10 ranks in the next snapshot.
@@ Results
Impact of Overall Engagement. The strongest factors associated with an increase in odds of staying near the top of the feed are the number of total comments, recent comments, and recent votes. We also found age (i.e., time since post creation) to be a strong factor associated with decreases in the odds of staying near the top. These are consistent across all rank intervals and are natural for an engagement-based popular feed.
We also observed that increasing the score on a post corresponds to a drastic decrease in the odds of staying on the feed. According to the values from each model, a substantial amount of variance remains unexplained. This suggests that unmodeled factors may be strongly correlated with score, contributing to the strong negative relationship observed in Table 2.
From our causal analysis, we found that increasing the number of comments increased a post’s tenure on r/popular in some cases. We found strong evidence that both the weak (TC-W) and the strong (TC-S) treatments had a significant effect on the odds of appearing within the top 10 (0.23% in ROPE), with a probability of 99.98% () that the effect was positive for the weak treatment (mean=17.59%, CI[9.53%, 25.86%]), and a probability of 99.88% () that the effect was positive for the strong treatment (mean=28.15%, CI[14.23%, 45.06%]). However, there was little evidence that either treatment significantly increased the odds of staying in the top 25 and top 50 (18.45% in ROPE). Thus, we were only able to causally link an increase in recent comments to tenure on the top 10.
Impact of Undesired Activity. According to Table 2, we found that increasing the proportion of undesired comments and the number of subscribers has a comparatively small but significant positive relationship with the odds that a post stays within the top 50 and top 25, with a stronger effect in the top 10. While these effects do not necessarily indicate that the ranking algorithm is intentionally designed to promote antisocial content to drive more user engagement, these results fail to rule it out. These relationships also hint at an interaction between undesired comments and the rank of a post, and that the relationship with undesired comments is strongest for posts at the top of the feed.
From our causal analysis, we found that increasing the proportion of undesired comments could increase a post’s tenure on r/popular in some limited cases. We found evidence that the strong version of the treatment (TU-S) had a significant effect on the odds of appearing in the top 10 (2.58% in ROPE), with a probability of 99.93% () that this effect is positive (mean=13.09%, CI[5.34%, 21.90%]). However, we found little evidence that the weak treatment (TU-W) had any significant effect on appearing within any rank threshold (64.73% in ROPE), or that the strong treatment had any significant effect on the odds of appearing the top 25 or top 50 (88.70% in ROPE).